OpenStudy (anonymous):

Find the vertex of the parabola whose equation is y = -2x^2 + 8x - 5. (2, 3) ?

OpenStudy (anonymous):

Vertex: -b/2a = -8/-4 = 2 -2(2)^2 + 8(2) -5 = 3 So, you are right, (2,3)